My Notes

Use the z-table to find the requested probabilities. Enter your answers to 4 decimal places.

(a)    P(z < 1.71) =

(b)    P(z ≥ −1.62) =


(c)    P(−2.49 < z < 2.98) =
Answer #1

(a)    P(z < 1.71) = 0.9564

(b)    P(z ≥ −1.62) = 1 - P(z < −1.62) = 1 - 0.0526 = 0.9474

(c)    P(−2.49 < z < 2.98) = P(z < 2.98) - P(z < −2.49 ) = 0.9986 -0.0064 = 0.9922

***please ask if you have any doubts.Happy to help you.Thank you.